A specimen of oil having an initial volume of 200 is subjected to a pressure increase of 1.6*10^6Pa and the volume is found to decrease by 0.15. What is the bulk modulus of the material?.
1
Expert's answer
2020-05-21T13:59:37-0400
"K=\frac{\Delta P V}{\Delta V}=\frac{(1.6\cdot 10^6)(200)}{0.15}=2.1\cdot 10^9\ Pa"